OpenStudy (anonymous):

P = a + b + c, solve for c "My answer is P/a+b=c" I am not sure please help and explain

OpenStudy (anonymous):

That is right

OpenStudy (anonymous):

The choices in my homework don't support it a. a. P – a + b = c b. P – a – b = c c. P / -a + b = c d. P / a - b = c

OpenStudy (anonymous):

It should be A.

OpenStudy (anonymous):

so it is p- not p/ ?

OpenStudy (anonymous):

Yes sorry I had got mixed up earlier it is a

OpenStudy (anonymous):

I still don't understand why P-a+b=c

OpenStudy (anonymous):

The answer is B First you subtract a from both sides then it leaves you with p-a=b+c then you subtract b from both sides and gives you p-a-b= c

OpenStudy (alexandervonhumboldt2):

a+b+c=p a+b+c-a-b=p-a-b c=p-a-b

OpenStudy (anonymous):

5x + 10y = -20, solve for y so in this case the answer will be a. y = -½ x - 2